OpenStudy (anonymous):

can someone help??? im about to have a nerd meltdown which expression shows X^2+11x+18?X^2+5x+6 answer choices f. x+3 g. x+9/x+3 h. 3 j. x+6/x+2

Directrix (directrix):

X^2+11x+18 ? X^2+5x+6 What is this question mark doing in the middle of the problem?

OpenStudy (anonymous):

nothing..........my bad?

Directrix (directrix):

Supposed to be divided by? Before a nerd meltdown, always check the problem.

OpenStudy (anonymous):

it is answer g we just simplify the terms and divide like terms out

OpenStudy (anonymous):

= x+9/x+3

Directrix (directrix):

@chemENGINEER --> How do you know that when we don't even know what the problem is?

OpenStudy (anonymous):

which expression shows x^2+11x+18/x^2+5x+6

Directrix (directrix):

@lilly21 --> PUHleez post the problem correctly before I have a badly written problem meltdown. Thanks.

Directrix (directrix):

Should be written as ( x^2+11x+18) / (x^2+5x+6) but I can understand it now. Thanks.

Directrix (directrix):

g. (x+9) / (x+3) ( x^2+11x+18) / (x^2+5x+6) = (x +9)(x+2) / (x +2)(x+3) = (x +9) / (x + 3)

OpenStudy (anonymous):

Thanks.